# Training Small Reasoning LLMs with Cognitive Preference Alignment

### v1

@@ Abstract
The reasoning capabilities of large language models (LLMs), such as OpenAI’s o1 and DeepSeek-R1, have seen substantial advancements through deep thinking. However, these enhancements come with significant resource demands, underscoring the need to explore strategies to train effective reasoning LLMs with far fewer parameters. A critical challenge is that smaller models have different capacities and cognitive trajectories than their larger counterparts. Hence, direct distillation of chain-of-thought (CoT) results from large LLMs to smaller ones can be sometimes ineffective and requires a huge amount of annotated data. In this paper, we introduce a novel framework called Critique-Rethink-Verify (CRV), designed for training smaller yet powerful reasoning LLMs. Our CRV framework consists of multiple LLM agents, each specializing in unique abilities: (i) critiquing the CoTs according to the cognitive capabilities of smaller models, (ii) rethinking and refining these CoTs based on the critiques, and (iii) verifying the correctness of the refined results. We further propose the cognitive preference optimization (CogPO) algorithm to enhance the reasoning abilities of smaller models by aligning thoughts of these models with their cognitive capacities. Comprehensive evaluations on challenging reasoning benchmarks demonstrate the efficacy of CRV and CogPO, which outperforms other training methods by a large margin.  [^1] Source codes, datasets and models will be released upon paper acceptance.
Training Small Reasoning LLMs with Cognitive Preference Alignment
Wenrui Cai1,2, Chengyu Wang2thanks:    Corresponding author., Junbing Yan2, Jun Huang2, Xiangzhong Fang1 1 Shanghai Jiao Tong University, Shanghai, China 2 Alibaba Cloud Computing, Hangzhou, China chengyu.wcy@alibaba-inc.com
@@ 1 Introduction
The remarkable progress in large language models (LLMs) has revolutionized NLP Zhao et al. (2023). More recently, leading models such as OpenAI’s o1222https://openai.com/o1/ and DeepSeek-R1 DeepSeek-AI (2025) leverage slow thinking mechanisms to solve complex tasks. Despite these capabilities, the scale of these models results in substantial computational demands. Consequently, there is a growing need to facilitate effective reasoning abilities with considerably fewer model parameters.
@@
Figure 1: A motivation example. Large models (right) apply vector-based algebraic abstraction to solve the problem, while small models (left) employ simple formulaic geometric decomposition. This trajectory mismatch underscores the inefficacy of direct CoT distillation across models with substantial capacity gaps.
@@
A straightforward approach to address this challenge is the direct distillation of Chain-of-Thought (CoT) outputs Wei et al. (2022a) or other deep thoughts (such as Tree-of-Thought Yao et al. (2023b)) from larger LLMs to smaller models. This technique is widely applied to improve the capacities of smaller LLMs Hsieh et al. (2023); Shridhar et al. (2022); Li et al. (2023); Yue et al. (2024). However, smaller LLMs [^3] In this work, we regard smaller LLMs as decoder-only language models typically with fewer than 10B parameters. inherently possess different capacities and cognitive trajectories when solving problems compared to their larger counterparts, as illustrated in Figure 1. Similar findings have also been presented in Li et al. (2022); Zhang et al. (2024); Hu et al. (2024); Li et al. (2024). This phenomenon indicates that straightforward distillation of CoTs from larger models can sometimes be ineffective and requires a huge amount of data due to the large capacity gap. Thus, a natural question arises: How can we improve the reasoning abilities of smaller LLMs to fit their own cognitive capacity?
@@
In this paper, we introduce “Critique-Rethink-Verify” (CRV), a novel approach to enhance the reasoning capabilities of smaller LLMs. CRV leverages multiple LLM agents, each with specialized functions and working in synergy. These functions include (i) critique the CoT output by considering the cognitive limits of smaller LLMs, (ii) rethink and refine these CoTs, integrating the feedback received from the previous module, and (iii) verify the accuracy and validity of the refined reasoning paths. Extending the direct preference optimization (DPO) technique Rafailov et al. (2023), we further propose the cognitive preference optimization (CogPO) algorithm to align reasoning abilities with the unique cognitive strengths of smaller LLMs. Ultimately, the reasoning performance of smaller LLMs can be significantly improved.
@@
In the experiments, the effectiveness of our approach is evaluated on several challenging reasoning benchmarks that are difficult for models with limited parameter sizess such as AIME 2024, MATH-500 Lightman et al. (2023), GPQA-Diamond Rein et al. (2023), and LiveCodeBench. The results show that the smaller LLMs trained using the CRV framework and the CogPO algorithm achieve better reasoning performance, outperforming other popular training methods. In summary, we make the following major contributions:
  • We present the CRV framework for training smaller yet powerful reasoning LLMs, leveraging multiple LLM agents where each agent specializes in unique tasks.
@@
  • We propose the CogPO algorithm that enhances the reasoning abilities of smaller models by aligning their reasoning processes with their cognitive capacities.
@@
  • Our evaluations on several challenging reasoning benchmarks demonstrate that combining CRV and CogPO significantly improves the reasoning performance of smaller models.
@@ 2.1 Prompting LLMs to Reason
Prompting strategies to improve reasoning in LLMs have become a critical focus. Initial studies showed that LLMs could perform basic reasoning tasks using meticulously crafted prompts, such as linguistic analysis Chen et al. (2021) and commonsense inference Latcinnik and Berant (2020); Shwartz et al. (2020). These single-step reasoning approaches laid the groundwork for advanced prompting techniques. To name a few, Chain-of-Thought (CoT) Wei et al. (2022b) prompting explicitly guides LLMs through step-by-step reasoning, enabling them to decompose complex problems into manageable intermediate reasoning steps. Tree-of-Thought (ToT) Yao et al. (2023a) prompting introduces a hierarchical structure to reasoning trajectories, allowing models to explore multiple solution paths. Furthermore, self-refine Shinn et al. (2023); Madaan et al. (2023) prompting incorporates verification checkpoints, where models validate intermediate results before advancing.
Figure 2: Overview of our proposed approach. The framework consists of two synergistic phases: (1) Generation of cognitively aligned data using the CRV system for direct distillation using SFT, and (2) CogPO training with dynamic $\beta$ adjustment, considering different levels of mini-tasks. Disclaimer: We use the Qwen logo as our backbones; however, any LLMs with sufficient capabilities can serve as the agents as well.
@@ 2.2 Reasoning LLMs
With the advancement of LLMs, model capabilities have steadily improved Chen and Varoquaux (2024); Bansal et al. (2024), even as their scale has been reduced. Models with approximately 7B to 14B parameters show remarkable performance, and their fine-tuning costs have become increasingly feasible. This has led to the emergence of specialized small models tailored for mathematical and code-related reasoning tasks such as Qwen-Math [^4] https://qwenlm.github.io/blog/qwen2.5-math/, Qwen-Coder [^5] https://qwenlm.github.io/blog/qwen2.5-coder-family/, and Macro-o1 Zhao et al. (2024).
Recent studies Shridhar et al. (2023); Yan et al. (2023); Liang et al. (2024); Yuan et al. (2024) have investigated fine-tuning methods to enhance the reasoning abilities of smaller models. By utilizing intermediate reasoning steps, LLMs can iteratively refine their outputs Jiang et al. (2024); Wang et al. (2024); Chen et al. (2025). This progressive methodology enables the development of compact models with reasoning capabilities that rival or exceed those of larger LLMs (over 100B), leading to the release of stronger reasoning models such as QwQ-32B [^6] https://qwenlm.github.io/blog/qwq-32b/ and DeepSeek-R1 DeepSeek-AI (2025). Our work studies the cognitive capacities of small models, beneficial for the community to train more powerful reasoning LLMs.
@@ 2.3 Alignment Training
To effectively train LLMs, a reinforcement learning stage is typically employed after the supervised fine-tuning (SFT) phase, which serves to improve the model’s alignment towards certain objectives. Reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) has shown effectiveness in aligning LLMs with human feedback. A potential drawback of RLHF is the explicit need for a reward model and the unstable RL training process. Direct preference optimization (DPO) Rafailov et al. (2023) trains LLMs based on chosen and rejected responses. Since the introduction of DPO, several approaches have been proposed to enhance its efficacy and efficiency. For example, CPO Xu et al. (2024) extends DPO to avoid generating adequate but not perfect machine translations. SimPO Meng et al. (2024) simplifies DPO by eliminating the reference model. KTO Ethayarajh et al. (2024) and NCA Chen et al. (2024) develop novel optimization goals that leverage unpaired data for model alignment. Furthermore, SPPO Wu et al. (2024b) employs on-policy sampling to generate preference data, outperforming off-policy DPO methods. In our work, we extend DPO to align reasoning abilities with the cognitive strengths of smaller LLMs.
@@ 3 Proposed Approach
We present the techniques of our CRV system and the CogPO training algorithm in detail.
@@ 3.1 Overall Framework
Our framework consists of two synergistic phases: (1) SFT with cognitively aligned data using the CRV system, and (2) CogPO with dynamic $\beta$ adjustment. As illustrated in Figure 2, the CRV system first refines training data tailored to the cognitive capacity of smaller LLMs, while CogPO further aligns reasoning preferences through difficulty-aware optimization. This design ensures the model initially acquires capacity-matched reasoning patterns, followed by the refinement of its decision boundaries through gap-sensitive learning.
@@ 3.2 The CRV System
The CRV system employs three LLM agents to construct the SFT dataset aligned with the cognitive limits of smaller models. The input to the system is an initial training set $\mathcal{D}_{\text{SFT}}=\{(x,y,r_{\text{orig}})\}$, where the three elements denote the problem, the answer, and the original reasoning process generated by any large reasoning model (e.g., DeepSeek-R1).
@@ 3.2.1 Critic
A LLM agent first evaluates the appropriateness of reasoning processes for the target small LLM (denoted as $\pi_{\text{base}}$). For each $(x,y,r_{\text{orig}})\in\mathcal{D}_{\text{SFT}}$, the Critic assesses $r_{\text{orig}}$ using the criteria of Cognitive Matching Degree, where the Critic checks whether the complexity of $r_{\text{orig}}$ aligns with the reasoning capacity of $\pi_{\text{base}}$. Specifically, the Critic classifies the reasoning processes into three subsets: i) $\mathcal{D}_{\text{easy}}$: cases where the reasoning process is overly simple, failing to guide $\pi_{\text{base}}$ in solving $x$; ii) $\mathcal{D}_{\text{med}}$: cases with adequate steps that enable successful problem solving; and iii) $\mathcal{D}_{\text{hard}}$: cases with redundant or excessively complex reasoning steps that exceed the comprehension of $\pi_{\text{base}}$.
Remarks. An intuitive approach is to use $\pi_{\text{base}}$ as the Critic. However, due to the small parameter size (e.g., 7B), certain CoTs exceeded $\pi_{\text{base}}$’s comprehension, rendering it incapable of reliable difficulty classification. Thus, we leverage the same LLM for the Rethinker and the Verifier (denoted as $\pi_{\text{large}}$) to serve as the Critic, while forcing it to “think” from the perspective of the small model $\pi_{\text{base}}$. Detailed analysis on the Critic choices and prompt templates are provided in experiments and appendix.
@@ 3.2.2 Rethinker
A LLM agent $\pi_{\text{large}}$ is tasked with rewriting reasoning processes to achieve the cognitive alignment. For each $(x,y,r_{\text{orig}})\in\mathcal{D}_{\text{easy}}$, the Rethinker expands $r_{\text{orig}}$ by adding missing logical steps while maintaining the correctness of $r_{\text{orig}}$, i.e.,
$r_{\text{easy*}}=\pi_{\text{large}}(x,y,r_{\text{orig}}).$ (1)
Similarly, for each $(x,y,r_{\text{orig}})\in\mathcal{D}_{\text{hard}}$, the Rethinker simplifies $r_{\text{orig}}$ by removing redundancies or using simpler method to solve the problem grounded in the correct answer:
$r_{\text{hard*}}=\pi_{\text{large}}(x,y,r_{\text{orig}}).$ (2)
Cases of the rewriting outcomes are shown in Table 8 and Table 9.

### v2

# Enhancing Reasoning Abilities of Small LLMs with Cognitive Alignment

@@ Abstract
The reasoning capabilities of large reasoning models (LRMs), such as OpenAI’s o1 and DeepSeek-R1, have seen substantial advancements through deep thinking. However, these enhancements come with significant resource demands, underscoring the need for training effective small reasoning models. A critical challenge is that small models possess different reasoning capacities and cognitive trajectories compared with their larger counterparts. Hence, directly distilling chain-of-thought (CoT) rationales from large LRMs to smaller ones can sometimes be ineffective and often requires a substantial amount of annotated data. In this paper, we first introduce a novel Critique-Rethink-Verify (CRV) system, designed for training smaller yet powerful LRMs. Our CRV system consists of multiple LLM agents, each specializing in unique tasks: (i) critiquing the CoT rationales according to the cognitive capabilities of smaller models, (ii) rethinking and refining these CoTs based on the critiques, and (iii) verifying the correctness of the refined results. Building on the CRV system, we further propose the Cognitive Preference Optimization (CogPO) algorithm to continuously enhance the reasoning abilities of smaller models by aligning their reasoning processes with their cognitive capacities. Comprehensive evaluations on challenging reasoning benchmarks demonstrate the efficacy of our CRV+CogPO framework, which outperforms other methods by a large margin. [^1] Source code is released in the EasyDistill toolkit Wang et al. (2025a). URL: https://github.com/modelscope/easydistill
Enhancing Reasoning Abilities of Small LLMs with Cognitive Alignment
Wenrui Cai1,2thanks: The work was conducted during the internship at Alibaba Cloud Computing., Chengyu Wang2thanks: Corresponding author., Junbing Yan2, Jun Huang2, Xiangzhong Fang1 1 Shanghai Jiao Tong University, Shanghai, China 2 Alibaba Cloud Computing, Hangzhou, China {cwrcwr,xzfang}@sjtu.edu.cn {chengyu.wcy,yanjunbing.yjb,huangjun.hj}@alibaba-inc.com
@@ 1 Introduction
The remarkable progress in language reasoning models (LRMs) has revolutionized NLP Zhao et al. (2023). Recently, leading models such as OpenAI’s o1222https://openai.com/o1/ and DeepSeek-R1 DeepSeek-AI (2025) have leveraged slow, deliberative thinking to solve complex tasks. Despite their impressive capabilities, the scale of these models results in substantial computational demands. Consequently, there is a growing need to train reasoning models with fewer parameters.
@@
Figure 1: A motivating example. Large models (right) apply vector-based algebraic abstraction to solve the problem, while small models (left) employ simple formulaic geometric decomposition. This trajectory mismatch underscores the inefficacy of direct CoT distillation across models with substantial capacity gaps.
@@
A straightforward approach to address this challenge is the direct distillation of Chain-of-Thought (CoT) rationales Wei et al. (2022a) or other deep thoughts (such as Tree-of-Thought Yao et al. (2023b)) from larger LRMs to smaller ones. This technique is widely applied to improve the capabilities of smaller LRMs Hsieh et al. (2023); Shridhar et al. (2022); Li et al. (2023); Yue et al. (2024). However, smaller models [^3] In this work, we regard smaller LLMs as decoder-only language models typically with fewer than 10B parameters. inherently exhibit different reasoning capacities and cognitive trajectories when solving problems compared to their larger counterparts, as illustrated in Figure 1. Similar findings have also been presented in Li et al. (2022); Zhang et al. (2024); Hu et al. (2024); Li et al. (2024). This phenomenon indicates that direct distillation of CoTs from larger models can sometimes be ineffective due to the large capacity gap. Thus, a natural question arises: How can we improve the reasoning abilities of smaller LRMs in a way that is aligned with their own cognitive capacity?
@@
In this paper, we introduce the “Critique-Rethink-Verify” (CRV) system, a novel approach to enhance the reasoning capabilities of smaller models. CRV leverages multiple LLM agents, each with specialized functions working in synergy: (i) critiquing CoT rationales by considering the cognitive limits of smaller LRMs, (ii) rethinking and refining these CoTs, integrating the feedback from previous critiques, and (iii) verifying the accuracy and validity of the refined reasoning paths. Extending the Direct Preference Optimization (DPO) technique Rafailov et al. (2023), we further propose the Cognitive Preference Optimization (CogPO) algorithm to align the reasoning process with the cognitive capacities of smaller LRMs, building upon the CRV system. Ultimately, the reasoning performance of smaller models can be improved effectively.
@@
We evaluate the effectiveness of our approach on several challenging reasoning benchmarks that are difficult for models with limited parameter sizes, such as AIME 2024, MATH-500 Lightman et al. (2023), GPQA-Diamond Rein et al. (2023), and LiveCodeBench. The results indicate that the small LRMs trained using the CRV+CogPO framework achieve outstanding reasoning performance.
In summary, our major contributions are:
  • We present the CRV system for training small yet powerful LRMs, based on multiple LLM agents specialized in unique tasks.
@@
  • We propose the CogPO algorithm, which continuously enhances the reasoning abilities of small models by aligning their reasoning processes with their cognitive capacities.
@@
  • Evaluations on challenging benchmarks demonstrate that the CRV+CogPO framework significantly improves the reasoning performance of small models, outperforming other popular training methods.
@@ 2 Related Work
In this section, we summarize the related work in the following three aspects.
@@ 2.1 Prompting LLMs to Reason
Prompting strategies to improve reasoning in large language models (LLMs) have become a critical focus. Initial studies showed that LLMs could perform basic reasoning tasks using carefully crafted prompts, such as linguistic analysis Chen et al. (2021) and commonsense inference Latcinnik and Berant (2020); Shwartz et al. (2020). Chain-of-Thought (CoT) Wei et al. (2022b) prompting explicitly guides LLMs through step-by-step reasoning, enabling them to decompose complex problems into manageable intermediate steps. Tree-of-Thought (ToT) Yao et al. (2023a) prompting introduces a hierarchical structure to reasoning trajectories, allowing models to explore multiple solution paths. Furthermore, self-refine Shinn et al. (2023); Madaan et al. (2023) prompting incorporates verification checkpoints, where models validate intermediate results before advancing.
Figure 2: Overview of our CRV+CogPO framework, consisting of two synergistic phases: (1) SFT training with cognitively aligned data generated by the CRV system, and (2) CogPO: dynamic $\beta$ adjustment preference optimization training using cognitive reasoning pairs with different quality gaps. Disclaimer: We use the Qwen logo as our backbone; however, any LLMs with sufficient capabilities can serve as agents as well.
@@ 2.2 Reasoning LLMs
With the advancement of LLMs, model capabilities have steadily improved Chen and Varoquaux (2024); Bansal et al. (2024). Models with approximately 7B to 14B parameters now show remarkable performance, and their fine-tuning costs have become increasingly feasible. This has led to the emergence of specialized small models tailored for mathematical and code-related reasoning tasks such as Qwen-Math [^4] https://qwenlm.github.io/blog/qwen2.5-math/, Qwen-Coder [^5] https://qwenlm.github.io/blog/qwen2.5-coder-family/, and Macro-o1 Zhao et al. (2024).
Recent studies Shridhar et al. (2023); Yan et al. (2023); Liang et al. (2024); Yuan et al. (2024); Wang et al. (2025b) have investigated fine-tuning methods to enhance the reasoning abilities of smaller models. By utilizing intermediate reasoning steps, LLMs can iteratively refine their outputs Jiang et al. (2024); Wang et al. (2024); Chen et al. (2025). This methodology facilitates the development of small reasoning models, particularly following the release of stronger reasoning models such as DeepSeek-R1 DeepSeek-AI (2025), QwQ-32B [^6] https://qwenlm.github.io/blog/qwq-32b/ and many others.
@@ 2.3 Alignment Training
To effectively train LLMs, a reinforcement learning stage is typically employed after the supervised fine-tuning (SFT) phase, serving to improve the model’s alignment toward specific objectives. Reinforcement learning from human feedback (RLHF) Ouyang et al. (2022) has shown effectiveness in aligning LLMs with human preferences. A potential drawback of RLHF is the explicit need for a reward model and the instability of RL training. Direct Preference Optimization (DPO) Rafailov et al. (2023) trains LLMs based on selected and rejected responses. Since the introduction of DPO, several approaches have been proposed to enhance its efficacy and efficiency. For example, CPO Xu et al. (2024) extends DPO to avoid generating adequate but not perfect machine translations. SimPO Meng et al. (2024) simplifies DPO by eliminating the reference model. KTO Ethayarajh et al. (2024) and NCA Chen et al. (2024) develop novel optimization objectives that leverage unpaired data for model alignment. Furthermore, SPPO Wu et al. (2024b) employs on-policy sampling to generate preference data, often outperforming off-policy DPO methods. In our work, we extend DPO to align reasoning abilities with the cognitive limits of small LLMs.
@@ 3 Proposed Approach
In this section, we present the techniques of our CRV system and the CogPO training algorithm.
@@ 3.1 Overall Framework
Our framework consists of two synergistic phases: (1) SFT with cognitively aligned data generated by the CRV system, and (2) CogPO with dynamic $\beta$ adjustment. As illustrated in Figure 2, the CRV system first refines data tailored to the cognitive capacity of smaller LRMs for SFT training, and CogPO further aligns reasoning preferences through suitability-aware optimization using pairs with different quality gaps. This design ensures that the model initially acquires capacity-matched reasoning patterns, followed by refinement of its decision boundaries through gap-sensitive learning. Here, the decision boundary refers to the model’s ability to judge whether the produced CoT is correct and aligned with its own cognitive capabilities, enabling it to successfully solve problems by following its CoT.
@@ 3.2 The CRV System
The CRV system employs LLM agents to construct an SFT dataset aligned with the cognitive limits of the smaller models to be trained. The input to the CRV system is an initial training set $\mathcal{D}_{\text{SFT}}=\{(x,y,r_{\text{orig}})\}$, where the three elements denote the problem, the correct answer, and the original reasoning process generated by any large LRM (e.g., DeepSeek-R1), which has been validated as correct. The following describes each agent in the CRV system.
@@ 3.2.1 Critic
An LLM agent first evaluates the appropriateness of reasoning processes for the target small LLM (denoted as $\pi_{\text{base}}$). For each $(x,y,r_{\text{orig}})\in\mathcal{D}_{\text{SFT}}$, the Critic assesses $r_{\text{orig}}$ using the criterion of Cognitive Matching Degree, checking whether the complexity and difficulty of $r_{\text{orig}}$ aligns with the cognitive capacity of $\pi_{\text{base}}$. Specifically, the Critic classifies the reasoning processes into three subsets: i) $\mathcal{D}_{\text{easy}}$ : $(x,y,r_{\text{easy}})$, cases where the reasoning process is overly terse, making it difficult for $\pi_{\text{base}}$ to follow; ii) $\mathcal{D}_{\text{med}}$ : $(x,y,r_{\text{med}})$, cases with appropriate steps that enable successful problem solving; and iii) $\mathcal{D}_{\text{hard}}$ : $(x,y,r_{\text{hard}})$, cases with overly redundant or excessively complex reasoning steps that exceed the comprehension of $\pi_{\text{base}}$, making it extremely likely to fail to guide $\pi_{\text{base}}$ in solving $x$.
Remarks. An intuitive approach would be to use $\pi_{\text{base}}$ itself as the Critic. However, due to its small parameter size (e.g., 7B), certain CoTs exceed $\pi_{\text{base}}$’s comprehension, rendering it incapable of reliable complexity classification. Thus, we leverage the same LLM for the Rethinker (denoted as $\pi_{\text{large}}$) to serve as the Critic, forcing it to “think” from the perspective of the small model $\pi_{\text{base}}$. A detailed analysis of choices for the Critic is provided in the Experiments (Section 4.3) and Appendix A.5.
Hypothesis Verification. To further verify that the complexity levels of CoTs are closely related to the cognitive capacities of reasoning models, we conduct an experiment in which we evaluate DeepSeek-R1-Distill-Qwen-1.5B/7B/32B on MATH500, collecting each model’s outputs. We employ the Critic to rate the level of each model’s CoT outputs; each CoT is evaluated three times, and the final rating is determined by majority vote. For each model, we quantify the distribution of these CoTs across different complexity levels in Table 1. As shown, DeepSeek-R1-Distill-Qwen-1.5B yields the largest number of simple CoTs, while DeepSeek-R1-Distill-Qwen-32B generates the greatest number of difficult CoTs.
These findings demonstrate that the complexity of CoTs escalates as the model size increases, suggesting that larger models possess higher reasoning and cognitive capacities. Consequently, overly terse or complex CoTs may not be suitable for training models with lower cognitive abilities. It is therefore essential to use CoTs that align with the model’s cognitive trajectory to improve its reasoning capabilities, a strategy akin to “teaching according to the student’s ability.”
Level/Model Size 1.5B 7B 32B
Easy 195 80 19
Medium 296 389 354
Hard 9 31 127
Table 1: Complexity distributions of CoTs generated by different sizes of DeepSeek-R1-Distill-Qwen models.
@@ 3.2.2 Rethinker
An LLM agent $\pi_{\text{large}}$ is tasked with rewriting reasoning processes to achieve cognitive alignment. For each $(x,y,r_{\text{easy}})\in\mathcal{D}_{\text{easy}}$, the Rethinker expands $r_{\text{easy}}$ by adding necessary steps for easier understanding, i.e., $r_{\text{easy*}}=\pi_{\text{large}}(x,y,r_{\text{easy}})$. Similarly, for each $(x,y,r_{\text{hard}})\in\mathcal{D}_{\text{hard}}$, the Rethinker simplifies $r_{\text{hard}}$ by removing redundancies or using simpler methods to solve the problem grounded in the correct answer: $r_{\text{hard*}}=\pi_{\text{large}}(x,y,r_{\text{hard}})$. Examples of the rewriting process of the Rethinker are shown in Tables 11 and 12.